OpenStudy (anonymous):

A county has constructed a conical building to store sand. The cone has a height of 195 ft and a diameter of 307 ft. Find the volume of this building to the nearest hundredth

OpenStudy (campbell_st):

The volume of a cone is \[V = \frac{1}{3} \pi r^2 h\] you know h = 195 and you can find the radius r by halving the diameter then substitute the values and calculate